OpenStudy (anonymous):

Find x3 - 2y2 - 3x3 + z4 if x = 3, y = 5, and z = -3.

OpenStudy (anonymous):

im not good in math

OpenStudy (camerondoherty):

Ok so we start Off by substituting z for -3 y for 5 x for 3 \[\Large x^3 - 2y^2 - 3x^3 + z^4\] \[\Large (3)^3 - 2(5)^2 - 3(3)^3 + (-3)^4\] So acording to pemdas we do exponents first \[\Large 3*3*3=27\] \[\Large 5^2 = 25\] \[\Large 3*3*3 = 27\] \[\Large -3*-3*-3*-3 = -81\] \[\Large 27 - 2(25) - 3(27) + -81\] Since The Next thing to do on pemdas is multiply we multiply \[\Large 2*25=50\] \[\Large 3*27=81\] \[\Large 27-50-81 - 81\] So can you solve that?

OpenStudy (camerondoherty):

So based on this the answer is -23

OpenStudy (camerondoherty):

and i meant +81 :/

OpenStudy (anonymous):

wait so its -23?

OpenStudy (anonymous):

Yes it is -23. Everything that Cameron did is correct.

OpenStudy (anonymous):

ok thanks cause I got -185